OpenStudy (anonymous):

*PLEASE HELP* 3x-y if x=-2 and y=-2

OpenStudy (anonymous):

Plug in -2 for x and -2 for y and evaluate: \[3x-y = 3(-2) - (-2)= -6+2=-4\]

OpenStudy (anonymous):

thanks